OpenStudy (anonymous):

how do I convert 4.5 into a fraction?

OpenStudy (diyadiya):

9/2

OpenStudy (anonymous):

multiply it by \[\frac{10}{10}\] and you get \[\frac{45}{10}\] cancel by 5 to get \[\frac{9}{2}\]